Ahmed bought a ticket of train for Rs 150. But his plan was postponed and he had to return the ticket some hours before the departure of the train. Ahmed was given back only Rs 120. What percentage was deducted by the ticket seller?
1
Expert's answer
2020-10-14T18:43:44-0400

Amount deducted by ticket seller is=150-120=30.

Hence the percentage deducted by the ticket seller is "\dfrac{30}{150}*100=20"

Answer is 20%